OpenStudy (anonymous):

In a certain rectangle, the length exceeds the width by 2m. If the length is increased by 4m and the width is decreased by 2m, the area of the new rectangle is the same as that of the original rectangle. Find the dimensions of the original rectangle. What is the width and what is the length?

OpenStudy (lilg132):

original dimensions width = x length = x + 2 new dimensions width = x - 2 l = x + 6 x(x + 2) = (x - 2) (x + 6) x^2 + 2x = x^2 + 6x - 2x - 12 x^2 + 2x = x^2 + 4x - 12 -2x = -12 x = 6

OpenStudy (lilg132):

original width = 6m original length = 8m new width = 4 new length = 12

OpenStudy (lilg132):

width * length = 48m

OpenStudy (lilg132):

dont forget to click good answer :)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (lilg132):

your welcome :)